A SEARCH FOR HYDROXYLAMINE (NH_{2}OH) TOWARDS IRC+10216, ORION-S, ORION(KL), SGRB2(N), SGRB2(OH), W51M AND W3(IRS5)

\underline{ROBIN L. PULLIAM}, ANTHONY J. REMIJAN, National Radio Astronomy Observatory, Charlottesville, VA 22903; JOANNA CORBY, Dept. of Astronomy, Dept. of Chemistry, University of Virginia and National Radio Astronomy Observatory, Charlottesville, VA 22903.

Over 50 molecular species containing nitrogen have been detected in the gas phase towards various astronomical environments. A number of chemical models have tried to reproduce the observed abundances of N-bearing species in the various environments of the ISM and CSM with limited success. Recently, there has been an interest in hydroxylamine (NH_{2}OH) because of its possible role in the formation of amino acids in space.^{a}^{,}^{b} Implications of amino acid formation in space and detections of possible precursor species would have a profound impact on our pursuit of understanding the prebiotic molecular origins of life. However, hydroxylamine has yet to be detected in the gas phase in the interstellar medium. A recent gas-grain model by Garrod et al. (2008)^{c} suggests that NH_{2}OH is formed through radical recombination on grain surfaces and predicts an abundance ratio range of NH_{2}OH/H_{2} between 1.2x10^{−6} - 3.5x10^{−7}; thus, suggesting that NH_{2}OH should be within the detectable limits of existing radio receivers. Here we present the search for hydroxylamine towards seven different sources: IRC+10216, Orion-S, Orion(KL), SgrB2(N), SgrB2(OH), W51M, and W3(IRS5) at 2mm wavelengths using archival data from the NRAO 12m telescope.^{d}^{,}^{e} Hydroxylamine is not confirmed toward any region and upper limits to the total column density were determined for each source. The implications of these upper limits of NH_{2}OH to both the gas phase and grain surface models will also be discussed.

aV. Blagojevic, S. Petrie, D. K. Bohme (2003) Mon. Not. R. Astron. Soc. 339, L7-L11

bJ. L. Snow, G. Orlova,V. Blagojevic, and D. K. Bohme (2007) J. Am. Chem. Soc. 129, 9910-9917

cR. T. Garrod, S. L. Widicus Weaver, and E. Herbst, (2008) ApJ 682, 283-302

dA. J. Remijan, D. P. Leigh,A. J. Markwick-Kemper, and B. E. Tuner (2008) arXiv:0802.2273 [astro-ph]

eThe data used are from survey data taken by Barry E. Turner between 1993-1995 at the NRAO 12m telescope

and available at: http://www.cv.nrao.edu/Turner2mmLineSurvey and astroph reference.